# Probabilistic Height Grid Terrain Mapping for Mining Shovels using LiDAR

### v1

@@ Abstract
This paper explores the question of creating and maintaining terrain maps in environments where the terrain changes. The specific example explored is the construction of terrain maps from 3D LiDAR measurements on an electric rope shovel. The approach extends the height grid representation of terrain to include a Hidden Markov Model in each cell, enabling confidence-based mapping of constantly changing terrain. There are inherent difficulties in this problem, including semantic labelling of the LiDAR measurements associated with machinery and determining the pose of the sensor. Solutions to both of these problems are explored. The significance of this work lies in the need for accurate terrain mapping to support autonomous machine operation.
@@ 1 Introduction
Here, we focus on generating a terrain map using a shovel-mounted 3D LiDAR. A terrain map is required to support real-time excavation sequencing, manage the interaction between the dipper and the dig face (knowing where and how to initiate the dig), and allow productivity monitoring.
The contributions of this work are to (i) provide an overview of a terrain mapping framework using a shovel-mounted LiDAR, (ii) detail a novel probabilistic approach to update a terrain map using LiDAR measurements, and (iii) demonstrate the approach using field datasets.
Generating a terrain map from a shovel-mounted sensor is a challenging task. A 3D LiDAR sensor provides a set of range measurements at specified heading and elevation angles. The sensor itself has range measurement uncertainty and requires accurate intrinsic and extrinsic calibration to provide accurate point cloud data, see Nouiraa et al. (2016); D’Adamo et al. (2018). The sensor’s pose is also required to locate point cloud measurements within a global frame. Errors in the sensor pose propagate to form incorrect beliefs about the environment, see Yin et al. (2024).
@@
Figure. 1 illustrates the shovel’s operating environment and the data captured from a shovel-mounted LiDAR. These measurements provide no semantic information and require interpretation to generate beliefs about the environment they represent. This is challenging as the point cloud captures measurements on the shovel, other machines, the terrain, and dust. The terrain is also dynamically changing, with material being removed and spilled in the workspace. Dusty environments and material spillage from the dig face and dipper lead to spurious measurements and an inaccurate representation of the environment, see Phillips et al. (2017). A robust approach is needed to map terrain in these challenging conditions.
@@ 2 Related work
Literature is rich with approaches to map static environments using point cloud data. Occupancy-based methods discretize the environment into voxels (or pixels in 2D) and label each voxel as occupied if a measurement is coincident with the voxel or free if the voxel intersects the beam’s path. However, these approaches are based on the assumption of a static environment and typically provide accurate results in indoor experiments only, see Thrun et al. (1998, 2000).
@@
A height grid representation is used by Souza et al. (2013) as an alternative volumetric representation to voxels. A 2D grid composed of cells represents the environment. Each cell has a height represented by a mean and standard deviation, updated using a probabilistic law derived from the Kalman filter formulation. As before, there is no handling of dynamic changes in the environment. Recent approaches to mapping a static environment, such as Removert (Kim and Kim, 2020), ERASOR (Lim et al., 2021), and DynamicFilter (Fan et al., 2022), aim to extract a static map of the environment using free space constraints to update the global map. The terrain mapping problem is different as the static environment changes with the extraction and spillage of material.
These ideas have been used to demonstrate terrain mapping from shovel-mounted sensors. Dunbabin and Corke (2006) generate a Digital Terrain Map (DTM) using a 2D laser scanner mounted to a 1/6 scale rope shovel to image the terrain when the shovel is swinging or in tuck. Using the DTM for autonomous excavation is listed as future work, and there are no details about handling scenarios with dust and other agents in the workspace. Hillier et al. (2015) generates DTMs by fusing RaDAR and LiDAR sensing technologies. The GNSS pose of the shovel is used to stitch successive point clouds, and haul trucks are labelled by segmenting points estimated to be associated with the truck by using its GNSS pose. The authors note the sensitivity of the approach to the uncertainty in the shovel and truck pose estimates. Bettens (2023) demonstrates a terrain mapping method using two excavator-mounted 3D LiDARs. The approach constructs a height grid representation of the dig face using a visibility-based approach where measurements in previously identified free space are disregarded. While this is generally not a problem in excavation scenarios, material added to the workspace (e.g., face collapses or the movement of material) cannot be modelled.
@@ 3 Methodology
Figure 2: The figure illustrates the conversion of the voxelized map (left) to a height grid (right) constructed using the voxelized measurements, $P_{\mathcal{S},k}^{{}^{\prime}}$, and the raycast result to find all the observed voxels, $P_{\mathcal{S},k}^{{}^{\prime}obs}$. For each observed cell, $c_{i}$, the maximum of consecutive measurements is used to update the global map, $M_{\mathcal{M},k}$. Measurements in $P_{\mathcal{S},k}^{{}^{\prime}}$ that are above observed free voxels in $P_{\mathcal{S},k}^{{}^{\prime}obs}$ are disregarded as they are likely to be noisy measurements from dust or missed in the semantic labelling process.
The following section details the proposed approach to represent and update a terrain map using a shovel-mounted 3D LiDAR. The novelty of the method is the ability to update dynamic changes quickly and efficiently to provide a real-time terrain map.
@@ 3.1 Map Representation
$\hat{\mathbf{x}}_{i,k}=\eta\mathbf{A}\mathbf{B}_{i,k}\hat{\mathbf{x}}_{i,k-1},$ (1)
@@
$\mathbf{A}=\begin{bmatrix}S_{11}&S_{12}&\dots&S_{1n}\\ S_{21}&S_{22}&\dots&S_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ S_{n1}&S_{n2}&\dots&S_{nn}\end{bmatrix}=\begin{bmatrix}0.99&\delta&\dots&\delta\\ \delta&0.99&\dots&\delta\\ \vdots&\vdots&\ddots&\vdots\\ \delta&\delta&\dots&0.99\end{bmatrix},$ (3)
@@
to ensure each row and column sum to one. The self-transition probability of each state is set to a large value of $0.99$ to allow for the accumulation of sufficient evidence before changing state. The transition probabilities are insensitive to small changes and do not severely impact the algorithm’s performance. Each observed cell is initialized in the state corresponding to its first measurement. Erroneous initialization is handled by allowing $m$ observations to incrementally update the global map before using it for subsequent planning stages. The global map may also be initialized by using a survey of the area to be mapped.
@@ 3.2 Map Update
The shovel is labelled by fitting known geometries of the house, boom, dipper assembly (handle, dipper, dipper door), and tracks at their estimated poses relative to the sensor and labelling all measurements close to these geometries. The boom and house are static relative to the sensor and are labelled using geometric constraints. The pose of the dipper assembly and tracks are estimated using pose estimation algorithms that provide sufficiently accurate performance, e.g. MSoE (Phillips and McAree, 2018) or PLuM (Bhandari et al., 2023). Methods to validate this semantic labelling, e.g. Phillips et al. (2016), are also useful in the labelling process. Machine data can be alternatively used if available. Other agents in the workspace, such as the trucks and clean-up equipment, can be identified by locating their GNSS pose relative to the sensor or using known geometry pose estimation algorithms such as MSoE and PLuM. Spurious measurements from dust are also labelled as terrain and handled in the process described below.
@@
The sensor’s pose is estimated using a GNSS solution or a LiDAR odometry method such as SiMpLE (Bhandari et al., 2024). The point cloud in the map frame, $P_{\mathcal{M},k}$, is discretized at a resolution of $\Delta$ to form voxelized scan $P_{\mathcal{M},k}^{{}^{\prime}}$. The measurement $z$-values from the scan in the map frame, $P_{\mathcal{M},k}$, are preserved for each voxel, $v$, to avoid quantization errors in the following steps.
A process is needed to convert the voxelized map to a height grid where each ($x,y$) coordinate is assigned a height. The process used assumes that terrain must extrude from the ground and each cell must only have a single height. This assumption does not hold for terrain with overhangs and is a known limitation of the selected height grid representation. Figure. 2 illustrates the process to aid the explanation below. First, a raycast is performed from the sensor’s pose to each voxelized measurement in $P_{\mathcal{M},k}^{{}^{\prime}}$ to find all the observed voxels, $P_{\mathcal{M},k}^{{}^{\prime}obs}$. All voxels in the LiDAR’s beam path from the sensor to the measurement are labelled as free and the measurement is labelled as occupied. For the $i$-th cell, $c_{i}\in P_{\mathcal{M},k}^{{}^{\prime}}$, all voxelized observations for the $z$-coordinate in $P_{\mathcal{M},k}^{{}^{\prime}obs}$ are saved in order of increasing height, $H_{i}=\{h_{0},h_{1},\ldots,h_{j}\}$. If $h_{0}$ is a free voxel, the cell is removed from the list of cells to be updated as terrain cannot exist above the observed space as assumed previously. This constraint avoids updating measurements corresponding to dust or measurements missed from the semantic labelling process. If $h_{0}$ is labelled as occupied, the height is incremented until a free space voxel or the last element of $H$ is reached. The unquantized height corresponding to this value, $h_{i}$, is assigned to cell $c_{i}$. This process is repeated for all cells to construct the height grid for this scan, $P_{\mathcal{M},k}^{h}=\{c_{i}\}$.
@@ 3.4 Summary and Limitations
Figure 3: A summary of the map update process. The new point cloud is first semantically labelled to estimate points associated with the terrain and then voxelized to find all the observed space. This information is used to convert to a height grid and compute the likelihood of each cell being in the $n$ states. The observation is fused with the existing map using the HMM filter.
@@ 4 Results
Figure 4: The plots display the net change in observed volume calculated for two datasets with a cell discretization size of $\Delta=0.25$ m and updated at a frequency of 1 Hz. The datasets are manually labelled to provide a reference of the shovel’s behaviour during the dataset collection. Red squares indicate the beginning of a dig and blue diamonds are the dumping of material in a haul truck. A net change of approximately $200\,\text{m}^{3}$ is observed between new truck loads in both datasets.
@@
The algorithm is configured to a map height range of $[h_{min},h_{max}]=[0,20]$ m, at a discretization size of $\Delta=0.25$ m, resulting in each cell having $81$ states to represent all possible heights within the desired range. The standard deviation, $\sigma$, used to construct the likelihood matrix, $\mathbf{B}_{i}$, is set to equal $\Delta$, and we find this to be an insensitive parameter. The PLuM pose estimation algorithm is used to semantically label the scans with the shovel’s pose and SiMpLE is used to estimate the sensor’s pose. The map is initialized as an empty height grid due to the absence of a workspace survey. Extraction rates are estimated after an initialization period of $m=1000$ scans ($<$ 2 minutes).
@@
Figure. 1 displays a typical scene with a model of the shovel and truck. The height grid is generated by recursively fusing observations into a global map. Figure. 4 displays the extraction rates for the two datasets. The plots indicate the net change in observed volume in the workspace calculated using (8). We find using 10 Hz data provides minimal change in the maps produced, and hence use a lower update frequency to support real-time mapping. All results presented provide map updates at 1 Hz using a CPU-based implementation.
@@
Example changes in terrain over a receding window of 1000 scans are illustrated in Fig. 5. Areas with material removed are coloured red, and areas with an addition of material are coloured blue. Material addition refers to material falling from the digface when disturbed by the dipper and spillage. Figure. 5 (left) shows two main areas excavated by the shovel in the receding window, with material addition on either side of the area likely caused by spillage from the dipper. The plot on the right shows material disturbed from the top of the digface and accumulated at the toe. Accounting for the addition and removal of material allows for a net volume change to be calculated.
@@ 5 Conclusions and Future Work
This paper introduces a novel approach to probabilistically fusing new observations into an existing map. Representing each cell using an HMM allows for robustness against spurious measurements from dust and the semantic labelling process. The results demonstrate real-time terrain mapping using field datasets from a CAT 7495 shovel.
Future work involves recording a survey before and after the test sequence to quantitatively evaluate the mapping accuracy. The terrain map estimates from the point cloud data can also be fused with material movement models as demonstrated by D’Adamo (2023) to allow for faster updates of changes in the terrain instead of relying on full observability of the scene. This is also expected to suppress noisy estimates and provide a more accurate representation of the environment.

### v2

@@ Abstract
This paper explores the question of creating and maintaining terrain maps in environments where the terrain changes. The specific example explored is the construction of terrain maps from 3D LiDAR measurements on an electric rope shovel. The approach extends the height grid representation of terrain to include a Hidden Markov Model in each cell, enabling confidence-based mapping of constantly changing terrain. There are inherent difficulties in this problem, including semantic labelling of the LiDAR measurements and determining the pose of the sensor. The significance of this work lies in the need for accurate terrain mapping to support autonomous machine operation.
@@ 1 Introduction
In this paper, we focus on generating a terrain map using a shovel-mounted 3D Light Detection and Ranging (LiDAR) sensor. A terrain map is required to support real-time excavation sequencing, manage the interaction between the dipper and the dig face (knowing where and how to initiate the dig), and allow productivity monitoring. The contributions of this work are to (i) provide an overview of a terrain mapping framework using a shovel-mounted LiDAR, (ii) detail a novel probabilistic approach to update a terrain map using LiDAR measurements, and (iii) demonstrate the approach using field datasets.
Generating a terrain map from a shovel-mounted sensor is a challenging task. A 3D LiDAR sensor provides a set of range measurements at specified heading and elevation angles. The sensor has range measurement uncertainty and requires intrinsic and extrinsic calibration to provide accurate point cloud data, see Nouiraa et al. (2016); D’Adamo et al. (2018). The sensor’s pose is also required to locate point cloud measurements within a global frame. Errors in the sensor pose propagate to form incorrect beliefs about the environment, see Yin et al. (2024).
@@
Figure. 1 illustrates the shovel’s operating environment and the data captured from a shovel-mounted LiDAR. These measurements provide no semantic information and require interpretation to generate beliefs about the environment they represent. This is challenging as the point cloud captures measurements on the shovel, other machines, the terrain, and dust. The terrain is also dynamically changing, with material being removed and spilled in the workspace. Dusty environments and material spillage from the dig face and dipper lead to spurious measurements and an unpredictable representation of the environment, see Phillips et al. (2017). A robust approach is needed to map terrain in these challenging conditions.
@@ 2 Related work
Literature is rich with approaches to map static environments using point cloud data. Occupancy-based methods discretize the environment into voxels (or pixels in 2D) and label each voxel as occupied if a measurement is coincident with the voxel, or free if the voxel intersects the beam’s path. However, these approaches are based on the assumption of a static environment and typically provide accurate results in indoor experiments only, see Thrun et al. (1998, 2000).
@@
A height grid representation is used by Souza et al. (2013) as an alternative volumetric representation to voxels. A 2D grid composed of cells represents the environment. Each cell has a height described by a mean and standard deviation, updated using a probabilistic law derived from the Kalman filter formulation. As before, there is no handling of dynamic changes in the environment. Recent approaches to mapping a static environment, such as Removert (Kim and Kim, 2020), ERASOR (Lim et al., 2021), and DynamicFilter (Fan et al., 2022), aim to extract a static map of the environment using free space constraints to update the global map. The terrain mapping problem is different as the static environment changes with the extraction and spillage of material.
These ideas have been used to demonstrate terrain mapping from shovel-mounted sensors. Dunbabin and Corke (2006) generate a Digital Terrain Map (DTM) using a 2D laser scanner mounted to a 1/6 scale rope shovel to image the terrain when the shovel is swinging or in tuck. Using the DTM for autonomous excavation is listed as future work, and there are no details about handling scenarios with dust and other agents in the workspace. Hillier et al. (2015) generates DTMs by fusing Radio Detection and Ranging (RaDAR) and LiDAR sensing technologies. The Global Navigation Satellite Systems (GNSS) pose of the shovel is used to stitch successive point clouds, and haul trucks are labelled by segmenting points estimated to be associated with the truck, located by its GNSS pose. The authors note the sensitivity of the approach to the uncertainty in the shovel and truck pose estimates. Bettens (2023) demonstrates a terrain mapping method using two excavator-mounted 3D LiDARs. The approach constructs a height grid representation of the dig face using a visibility-based approach where measurements in previously identified free space are disregarded. While this is generally not a problem in excavation scenarios, material added to the workspace (e.g., face collapses or the movement of material) cannot be modelled.
@@ 3 Methodology
Figure 2: The conversion of the voxelized map (left) to a height grid (right) constructed using the voxelized measurements, $P_{\mathcal{M},k}^{{}^{\prime}}$, and the raycast result to find all the observed voxels, $P_{\mathcal{M},k}^{{}^{\prime}obs}$. For each observed cell, $c_{i}$, the maximum of consecutive measurements is used to update the global map, $M_{\mathcal{M},k}$. Measurements in $P_{\mathcal{M},k}^{{}^{\prime}}$ that are above observed free voxels in $P_{\mathcal{M},k}^{{}^{\prime}obs}$ are disregarded as they are likely to be noisy measurements from dust or missed in the semantic labelling process.
The following section details the proposed approach to represent and update a terrain map using a shovel-mounted 3D LiDAR. The novelty of the method is the ability to update dynamic changes quickly and efficiently to provide a real-time terrain map. The mathematical foundation of the general HMM framework applied to this problem is adapted from James et al. (2024).
@@ 3.1 Map Representation
$\hat{\mathbf{x}}_{i,k}=\eta_{i,k}\mathbf{B}_{i,k}\mathbf{A}\hat{\mathbf{x}}_{i,k-1},$ (1)
@@
$\mathbf{A}=\begin{bmatrix}a_{11}&a_{12}&\dots&a_{1n}\\ a_{21}&a_{22}&\dots&a_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n2}&\dots&a_{nn}\end{bmatrix}=\begin{bmatrix}0.99&\delta&\dots&\delta\\ \delta&0.99&\dots&\delta\\ \vdots&\vdots&\ddots&\vdots\\ \delta&\delta&\dots&0.99\end{bmatrix},$ (3)
@@
to ensure each row and column sum to one. The self-transition probability of each state is set to a large value of $0.99$ to allow for the accumulation of sufficient evidence before changing state. The off-diagonal terms are set to low values of $\delta$ to allow for transitions from any state. Each observed cell is initialized in the state corresponding to its first measurement. Erroneous initialization is handled by allowing $m$ observations to incrementally update the global map before using it for subsequent planning stages. The global map may also be initialized by using a survey of the area to be mapped.
@@ 3.2 Map Update
The shovel is labelled by fitting known geometries of the house, boom, dipper assembly (handle, dipper, dipper door), and tracks at their estimated poses relative to the sensor and labelling all measurements close to these geometries. The boom and house are static relative to the sensor and are labelled using geometric constraints. The pose of the dipper assembly and tracks are estimated using pose estimation algorithms that provide sufficiently accurate performance, e.g. MSoE (Phillips and McAree, 2018) or PLuM (Bhandari et al., 2023). Methods to validate this semantic labelling, e.g. Phillips et al. (2016), are also useful in this process. Machine data can be alternatively used if available. Other agents in the workspace, such as the trucks and clean-up equipment, can be identified by locating their GNSS pose relative to the sensor or using the geometry pose estimation algorithms mentioned previously. Spurious measurements from dust are also labelled as terrain and handled in the process described below.
@@
The sensor’s pose is estimated using a Kalman filter to fuse measurements from GNSS and LiDAR odometry approaches such as SiMpLE (Bhandari et al., 2024). The point cloud in the map frame, $P_{\mathcal{M},k}$, is discretized at a resolution of $\Delta$ to form voxelized scan $P_{\mathcal{M},k}^{{}^{\prime}}$. The measurement $z$-values from the scan in the map frame, $P_{\mathcal{M},k}$, are preserved for each voxel, $v$, to avoid quantization errors in the following steps.
A process is needed to convert the voxelized map to a height grid where each ($x,y$) coordinate is assigned a height. The process used assumes that terrain must extrude from the ground and each cell must only have a single height. This assumption does not hold for terrain with overhangs and is a known limitation of the selected height grid representation (Bettens, 2023). Figure. 2 illustrates the process to aid the explanation below. First, a raycast is performed from the sensor’s pose to each voxelized measurement in $P_{\mathcal{M},k}^{{}^{\prime}}$ to find all the observed voxels, $P_{\mathcal{M},k}^{{}^{\prime}obs}$. All voxels in the LiDAR’s beam path from the sensor to the measurement are labelled as free and the measurement is labelled as occupied. For the $i$-th cell, $c_{i}\in P_{\mathcal{M},k}^{{}^{\prime}}$, all voxelized observations for the $z$-coordinate in $P_{\mathcal{M},k}^{{}^{\prime}obs}$ are saved in order of increasing height, $H_{i}=\{h_{0},h_{1},\ldots,h_{j}\}$. If $h_{0}$ is a free voxel, the cell is removed from the list of cells to be updated as terrain cannot exist above the observed space as assumed previously. This constraint avoids updating measurements corresponding to dust or measurements missed from the semantic labelling process. If $h_{0}$ is labelled as occupied, the height is incremented until a free space voxel or the last element of $H$ is reached. The unquantized height corresponding to this value, $h_{i}$, is assigned to cell $c_{i}$. This process is repeated for all cells to construct the height grid for this scan, $P_{\mathcal{M},k}^{h}=\{c_{i}\}$.
@@ 3.4 Summary and Limitations
Figure 3: A summary of the map update process. The point cloud is first semantically labelled to estimate points associated with the terrain and then voxelized to find the observed space. This information is used to convert to a height grid and compute the likelihood of each cell being in the $n$ states. The observation is fused with the existing map using the HMM filter.
@@ 4 Results
Figure 4: The net change in observed volume calculated for two datasets. The datasets are manually labelled to provide a reference of the shovel’s behaviour during the dataset collection. Red squares indicate the beginning of a dig and blue diamonds are the dumping of material in a haul truck.
@@
The algorithm is configured to a map height range of $[h_{min},h_{max}]=[0,20]$ m, at a discretization size of $\Delta=0.25$ m, resulting in each cell having $81$ states to represent all possible heights within the desired range. The standard deviation, $\sigma$, used to construct the likelihood matrix, $\mathbf{B}_{i}$, is set to equal $\Delta$. The cell size is also set to 0.25m. The map is initialized as an empty height grid due to the absence of a workspace survey. Extraction rates are estimated after an initialization period of $m=1000$ scans ($<$ 2 minutes).
@@
Figure. 1 displays a typical scene with a model of the shovel and truck. The height grid is generated by recursively fusing observations into a global map. Figure. 4 displays the extraction rates for the two datasets. The plots indicate the net change in observed volume in the workspace calculated using (8). We find using 10 Hz data provides a minimal change in the maps produced, and hence use a lower update frequency of 1 Hz to reduce CPU load. The proposed algorithm is capable of providing map updates at 10 Hz using multithreading on a 12th Gen Intel Core i5-1240P×16 with 16GB RAM.
@@
Example changes in terrain over a receding window of 1000 scans are illustrated in Fig. 5. Areas with material removed are coloured red, and areas with an addition of material are coloured blue. Material addition refers to material falling from the digface when disturbed by the dipper, and spillage. Figure. 5 (left) shows two main areas excavated by the shovel in the receding window, with material addition on either side of the area likely caused by spillage from the dipper. The plot on the right shows material disturbed from the top of the digface and accumulated at the toe. Accounting for the addition and removal of material allows for a net volume change to be calculated.
@@ 5 Conclusions and Future Work
This paper introduces a novel approach to probabilistically fuse new observations into an existing map. Representing each cell using an HMM allows for robustness against spurious measurements from dust and the semantic labelling process. The results demonstrate real-time terrain mapping using field datasets from a CAT 7495 shovel.
Future work involves recording a survey before and after the test sequence to evaluate the mapping accuracy and provide a quantitative comparison with existing mapping frameworks. Material movement models can also be used to allow for map updates in areas unobserved by the LiDAR or provide a secondary measurement source to fuse with point cloud data as demonstrated by D’Adamo (2023). This is important when applying the same mapping framework to other mining agents such as dozers, as it may be difficult to capture terrain transformed by the blade in the point cloud data directly.